# Riesz Representer Fitting under Bregman Divergence: A Unified Framework for Debiased Machine Learning

### v1

@@ Appendix A Related Work
A parallel line of work is density ratio estimation, which has been extensively studied in machine learning. Density ratios are used not only for semiparametric analysis but also in tasks such as anomaly detection (Hido et al., 2008). As discussed in this study and in existing work such as Uehara et al. (2020) and Lin et al. (2023), density ratio estimation is closely related to propensity score estimation. In particular, this study shows that the formulations of Riesz regression and LSIF in density ratio estimation are essentially the same. While Riesz regression applies to more general problems, the LSIF literature provides a range of theoretical and empirical results. One important extension is to generalize LSIF via Bregman divergence minimization (Sugiyama, 2006), and our study is strongly inspired by that work.
Appendix B Preliminary
@@ B.1 Rademacher complexity
B.2 Local Rademacher complexity bound
@@
Proposition B.1 (From Theorem 2.1 in Bartlett et al. (2005)).
@@ B.3 Bracketing entropy
B.4 Talagrand’s concentration inequality
@@
Proposition B.2 (Talagrand’s Lemma).
Let $\phi\colon{\mathbb{R}}\to{\mathbb{R}}$ be a Lipschitz continuous function with a Lipschitz constant $L>0$. Then, it holds that
@@
Appendix C Basic inequalities
C.1 Strong convexity
Lemma C.1 ($L_{2}$ distance bound from Lemma 4 in Kato & Teshima (2021)).
If $\inf_{h\in(-\infty),\infty}g^{\prime\prime}(\alpha)>0$, then there exists $\mu>0$ such that for all $h\in{\mathcal{H}}$,
$\|h-\alpha_{0}\|^{2}_{2}\leq\frac{2}{\mu}\Big(\text{BD}_{g}(\alpha)-\text{BD}_{g}(\alpha_{0})\Big)$
@@ C.1 Strong convexity
From the strong convexity and Lemma C.1, we have
@@
$\widehat{\alpha}\coloneqq\operatorname*{arg\,min}_{h\in{\mathcal{H}}}\widehat{\text{BD}}_{g}(\alpha)+\lambda J(\alpha),$
@@ C.3 Risk bound
Then, from Proposition C.2, we have
@@
Proof of Proposition C.2 is trivial. We prove Proposition C.3 below.
@@
From the strong convexity and Lemma C.1, we have
@@
From Proposition C.2, we have
@@
Appendix D Proof of Theorem 5.1
@@
in Proposition C.3. We can bound this term by using the empirical-process arguments.
Note that since $\alpha_{0}\in{\mathcal{H}}$, it holds that $\alpha^{*}=\alpha_{0}$, which implies that
D.1 Preliminary
@@
Definition D.1 (Derived function class and bracketing entropy (from Definition 4 in Kato & Teshima (2021))).
Given a real-valued function class $\mathcal{F}$}, define $\ell\circ{\mathcal{F}}:=\{\ell\circ f\colon f\in{\mathcal{F}}\}$. By extension, we define $I:\ell\circ{\mathcal{H}}\to[1,\infty)$ by $I(\ell\circ h)=I(\alpha)$ and $\ell\circ{\mathcal{H}}_{M}:=\{\ell\circ h:h\in{\mathcal{H}}_{M}\}$. Note that, as a result, $\ell\circ{\mathcal{H}}_{M}$ coincides with $\{\ell\circ h\in\ell\circ{\mathcal{H}}:I(\ell\circ h)\leq M\}$.
Proposition D.1.
@@
$H_{B}\big(\delta,\ell\circ{\mathcal{H}},\|\cdot\|_{L_{2}(P_{0})}\big)$ $\leq\frac{(s+1)(2v)^{\gamma}}{\gamma}\left(\frac{M}{\delta}\right)^{\gamma}.$
@@
$\sup_{\ell\circ h\in\ell\circ{\mathcal{H}}_{M}}\|\ell\circ h-\ell\circ\alpha^{*}\|_{L_{2}(P_{0})}$ $\leq c_{0}vM,$
$\sup_{\stackrel{{\scriptstyle\ell\circ h\in\ell\circ{\mathcal{H}}_{M}}}{{\|\ell\circ h-\ell\circ\alpha^{*}\|_{L_{2}(P_{0})}\leq\delta}}}\|\ell\circ h-\ell\circ\alpha^{*}\|_{\infty}$ $\leq c_{0}vM,\quad\text{for all }\delta>0.$
Proposition D.2 (Lemma 5.13 in van de Geer (2000), Proposition 1 in Kanamori et al. (2012)).
@@
D.2 Upper bound using the empirical-process arguments
From Propositions D.1D.3, we obtain the following result.
Proposition D.4.
@@ D.2 Upper bound using the empirical-process arguments
$=O_{p}\left(\max\left\{\frac{\|\widehat{\alpha}-\alpha^{*}\|_{L^{2}(P_{0})}^{1-\gamma/2}\left(1+\left\|\widehat{\alpha}\right\|_{{\mathcal{H}}}\right)^{1+\gamma/2}}{\sqrt{n}},\frac{\left(1+\left\|\widehat{\alpha}\right\|_{{\mathcal{H}}}\right)^{2}}{n^{2/(2+\gamma)}}\right\}\right),$
@@
From Proposition C.3 and $\alpha_{0}\in\cal\alpha^{\text{RKHS}}$, we have
$\left\|\widehat{\alpha}(X)-\alpha_{0}(X)\right\|^{2}_{L_{2}(P_{0})}+\lambda\|\widehat{\alpha}\|^{2}_{{\mathcal{H}}}$
$\leq\left({\mathbb{E}}-\widehat{{\mathbb{E}}}\right)\left[L(\widehat{\alpha},D,X)-L(\alpha_{0},D,X)\right]+\lambda\|f_{0}\|^{2}_{{\mathcal{H}}}.$
From Proposition D.4, we have
$\left\|\widehat{\alpha}(X)-\alpha_{0}(X)\right\|^{2}_{L_{2}(P_{0})}+\lambda\|\widehat{f}\|^{2}_{{\mathcal{H}}}$
$=O_{p}\left(\max\left\{\frac{\|\widehat{\alpha}-\alpha_{0}\|_{L^{2}(P_{0})}^{1-\gamma/2}\left(1+\left\|\widehat{f}\right\|_{{\mathcal{H}}}\right)^{1+\gamma/2}}{\sqrt{n}},\frac{\left(1+\left\|\widehat{\alpha}\right\|_{{\mathcal{H}}}\right)^{2}}{n^{2/(2+\gamma)}}\right\}\right)+\lambda\|r_{0}\|^{2}_{{\mathcal{H}}}.$
@@
$\left\|\widehat{\alpha}(X)-\alpha_{0}(X)\right\|^{2}_{L_{2}(P_{0})}+\lambda\|\widehat{f}\|^{2}_{{\mathcal{H}}}=O_{p}(\lambda),$ (7)
$\left\|\widehat{\alpha}(X)-\alpha_{0}(X)\right\|^{2}_{L_{2}(P_{0})}+\lambda\|\widehat{f}\|^{2}_{{\mathcal{H}}}=O_{p}\left(\frac{\|\widehat{f}-f_{0}|_{L^{2}(P_{0})}^{1-\gamma/2}\left(1+\left\|\widehat{f}\right\|_{{\mathcal{H}}}\right)^{1+\gamma/2}}{\sqrt{n}}\right),$ (8)
$\left\|\widehat{\alpha}(X)-\alpha_{0}(X)\right\|^{2}_{L_{2}(P_{0})}+\lambda\|\widehat{f}\|^{2}_{{\mathcal{H}}}=O_{p}\left(\frac{\left(1+\left\|\widehat{f}\right\|_{{\mathcal{H}}}\right)^{2}}{n^{2/(2+\gamma)}}\right).$ (9)
@@
$\lambda\|\widehat{f}\|^{2}_{{\mathcal{H}}}=O_{p}(\lambda).$
Therefore, we have $\left\|\widehat{\alpha}(X)-\alpha_{0}(X)\right\|_{P_{0}}=O_{p}(\lambda^{1/2})$ and $\|\widehat{r}\|_{{\mathcal{H}}}=O_{p}(1)$.
@@
$\lambda\|\widehat{f}\|^{2}_{{\mathcal{H}}}=O_{p}\left(\frac{\|\widehat{f}-f_{0})\|_{L^{2}(P_{0})}^{1-\gamma/2}\left(1+\left\|\widehat{f}\right\|_{{\mathcal{F}}}\right)^{1+\gamma/2}}{\sqrt{n}}\right).$
@@
$\lambda\|\widehat{f}\|^{2}_{{\mathcal{H}}}$ $=O_{p}\left(\frac{\|\widehat{f}-f_{0})\|_{L^{2}(P_{0})}^{1-\gamma/2}\left(1+\left\|\widehat{f}\right\|_{{\mathcal{F}}}\right)^{1+\gamma/2}}{\sqrt{n}}\right)$
@@
$\|\widehat{f}\|_{{\mathcal{H}}}$ $=O_{p}\left(\frac{\left(1+\left\|\widehat{f}\right\|_{{\mathcal{F}}}\right)^{2}}{\lambda^{1/2}n^{2/(2+\gamma)}}\right)=o_{p}(1).$
@@
$\lambda\|\widehat{f}\|^{2}_{{\mathcal{H}}}=O_{p}\left(\frac{\left(1+\left\|\widehat{f}\right\|_{{\mathcal{F}}}\right)^{2}}{n^{2/(2+\gamma)}}\right).$
As well as the argument in (8), we have $\|\widehat{r}\|_{{\mathcal{H}}}=o_{p}(1)$. Therefore, we have
@@
Appendix E Proof of Theorem 5.2
@@
E.1 Proof of Lemma E.1
We prove Lemma E.1 by bounding (6) in Proposition C.3.
@@
Lemma E.2 (Corresponding to (26) in Zheng et al. (2022)).
Suppose that the conditions in Lemma E.1 hold. Then, for any $z>0$, with probability $1-\exp(-z)$ it holds that
@@
Lemma E.3 (Corresponding to (29) in Zheng et al. (2022)).
Suppose that the conditions in Lemma E.1 hold. If there exists $u_{0}>0$ such that
@@
Proposition E.4 (From (32) in Zheng et al. (2022)).
@@
Proposition E.5 (Corresponding to (36) in Step 3 of Zheng et al. (2022)).
Suppose that the conditions in Lemma E.1 hold. Then, there exists a universal constant $C>0$ such that
@@
Proposition E.6 (Upper bound of the Rademacher complexity).
Suppose that the conditions in Lemma E.1 hold. If $n\geq\text{Pdim}({\mathcal{F}}^{\text{FNN}})$, $u_{0}\geq 1/n$, and $n\geq(2eM)^{2}$, we have
@@
Then, we prove Lemma E.1 as follows:
Proof of Lemma E.1.
@@
then from (6) and Lemmas E.2 and E.3, for every $z>0$, there exists a constant $C>0$ independent $n$ such that
@@
If there exists $j\leq\ell$ such that $\widehat{f}\in\overline{{\mathcal{S}}}^{\text{FNN}}_{j}$, then from (E.1), for every $z>0$, with probability at least $1-8\exp(-z)$, there exists a constant $C>0$ independent of $n$ such that
@@
Here, to obtain (14), we used $\overline{u}\geq\max\left\{\sqrt{\log(n)/n},4\sqrt{3}M\sqrt{z/n},u^{\dagger}\right\}$, (E.1), (12), and (13).
From Proposition E.5, it holds that
@@
E.2 Proof of Lemma E.2
From Proposition B.1, we have
@@
This is a direct consequence of Proposition B.1. Note that $\alpha^{*}$ and $\alpha_{0}$ are fixed, and it is enough to apply the standard law of large numbers; that is, we do not have to consider the uniform law of large numbers. However, we can still apply Proposition B.1, which is a general than the standard law of large numbers, with ignoring the Rademacher complexity part.
@@
Then, from Proposition B.1, for every $z>0$, with probability at least $1-\exp(-z)$, it holds that
@@
Appendix F Overfitting Problems
Non-negative Bregman divergence
Telescoping density ratio estimation
$\frac{p_{0}(x)}{p_{m}(x)}=\prod_{k=0}^{m-1}\frac{p_{k}(x)}{p_{k+1}(x)}.$

### v2

@@ A.2 ATE Estimation
Covariate balancing.
Covariate balancing is a popular approach for propensity score or balancing weight estimation. The propensity score is a balancing score, and based on this property, existing studies propose estimating the propensity score or the weights so that weighted covariate moments match between treated and control groups. Imai & Ratkovic (2013b) proposes estimating the propensity score by matching first moments, and Hazlett (2020) extends this idea to higher-moment matching by mapping covariates into a high-dimensional space via basis functions. On the other hand, methods that do not directly specify a propensity score model have also been proposed. Such methods are called empirical balancing and include entropy balancing (Hainmueller, 2012) and stable weights (Zubizarreta, 2015). These two approaches may appear different, but Zhao (2019) and Bruns-Smith et al. (2025) show that they are essentially equivalent through a duality relationship.
A.3 Density Ratio Estimation
A parallel line of work is density ratio estimation, which has been extensively studied in machine learning. We refer to the ratio between two densities as the density ratio. Density ratio estimation has attracted considerable attention as an essential task in various machine learning problems, such as regression under covariate shift (Shimodaira, 2000; Kato et al., 2024a), learning with noisy labels (Liu & Tao, 2016; Fang et al., 2020), anomaly detection (Smola et al., 2009; Hido et al., 2008; Abe & Sugiyama, 2019), two-sample testing (Keziou & Leoni-Aubin, 2005; Kanamori et al., 2010; Sugiyama et al., 2011a), change-point detection (Kawahara & Sugiyama, 2009), causal inference (Uehara et al., 2020), and recommendation systems (Togashi et al., 2021). While the density ratio can be estimated by separately estimating each density, such an approach may magnify estimation errors by compounding two independent estimations. To address this issue, end-to-end, direct density ratio estimation methods have been studied, including moment matching (Huang et al., 2007; Gretton et al., 2009), probabilistic classification (Qin, 1998; Cheng & Chu, 2004), density matching (Nguyen et al., 2010), density ratio fitting (Kanamori et al., 2009), and PU learning (Kato et al., 2019). It is also known that when complicated models such as neural networks are used for this task, the loss function can diverge in finite samples (Kiryo et al., 2017). Therefore, density ratio estimation methods with neural networks have been investigated (Kato & Teshima, 2021; Rhodes et al., 2020).
@@ A.3 Density Ratio Estimation
Density ratios are used not only for semiparametric analysis but also in tasks such as anomaly detection (Hido et al., 2008). As discussed in this study and in existing work such as Uehara et al. (2020) and Lin et al. (2023), density ratio estimation is closely related to propensity score estimation. In particular, this study shows that the formulations of Riesz regression and LSIF in density ratio estimation are essentially the same. While Riesz regression applies to more general problems, the LSIF literature provides a range of theoretical and empirical results. One important extension is to generalize LSIF via Bregman divergence minimization (Sugiyama, 2006), and this study is strongly inspired by that work.
In causal inference, Uehara et al. (2020) investigates efficient ATE estimation and policy learning under covariate shift. Kato et al. (2024b) applies this approach to adaptive experimental design, and Kato et al. (2025) extends the framework to a PU learning setup.
Appendix B Proof of the Automatic Covariate Balancing Property
B.1 Constrained Optimization Problem
where we replace $\alpha_{0}$ with $\widehat{\alpha}$. Then, it holds that
$\min_{\bm{\alpha}\in{\mathbb{R}}^{n}}$ $\ \ \ \frac{1}{n}\sum^{n}_{i=1}g(\alpha_{i})$
subjectto $\ \ \ \left|\frac{1}{n}\sum^{n}_{i=1}\Big(\alpha_{i}\partial g\big(\alpha(X_{i})\big)-m(W_{i},(\partial g)\circ\phi_{j})\Big)\right|\leq\lambda\qquad j=1,\dots,p.$
Linearity for the basis functions.
Next, we consider the case where
$\partial g\big(\alpha_{{\bm{\beta}}}(X_{i})\big)=\sum^{p}_{j=1}{\bm{\beta}}_{j}\widetilde{g}\left(X_{i},\phi_{j}(X_{i})\right).$
We consider the following constrained optimization problem:
$\min_{\bm{\alpha}\in{\mathbb{R}}^{n}}$ $\ \ \ \frac{1}{n}\sum^{n}_{i=1}g(\alpha_{i})$
subjectto $\ \ \ \left|\frac{1}{n}\sum^{n}_{i=1}\Big(\alpha_{i}\widetilde{g}\left(X_{i},\phi_{j}(X_{i})\right)-m(W_{i},(\partial g)\circ\phi_{j})\Big)\right|\leq\lambda\qquad j=1,\dots,p.$
B.2 Dual Formulation
Using Lagrange multipliers ${\bm{\beta}}_{j}\in{\mathbb{R}}$ ($j=1,2,\dots,p$), the constrained optimization problem can be written as
$\min_{{\bm{\alpha}}\in{\mathcal{A}}^{n}}\sup_{{\bm{\beta}}\in{\mathbb{R}}^{p}}$ $\ \ \ \left\{\frac{1}{n}\sum^{n}_{i=1}g(\alpha_{i})+\sum^{p}_{j=1}{\bm{\beta}}_{j}\left(\frac{1}{n}\sum^{n}_{i=1}\Big(m(W_{i},(\partial g)\circ\phi_{j})-\alpha_{i}\widetilde{g}\left(X_{i},\phi_{j}(X_{i})\right)\Big)-\operatorname{sign}({\bm{\beta}}_{j})\lambda\right)\right\}.$
The dual problem of the above constrained problem is
$\max_{{\bm{\beta}}\in{\mathbb{R}}^{p}}\inf_{{\bm{\alpha}}\in{\mathcal{A}}^{n}}$ $\ \ \ \left\{\frac{1}{n}\sum^{n}_{i=1}g(\alpha_{i})+\sum^{p}_{j=1}\left({\bm{\beta}}_{j}\frac{1}{n}\sum^{n}_{i=1}\Big(m(W_{i},(\partial g)\circ\phi_{j})-\alpha_{i}\widetilde{g}\left(X_{i},\phi_{j}(X_{i})\right)\Big)-|{\bm{\beta}}_{j}|\lambda\right)\right\}.$
$\max_{{\bm{\beta}}\in{\mathbb{R}}^{p}}\inf_{{\bm{\alpha}}\in{\mathcal{A}}^{n}}$ $\ \ \ \frac{1}{n}\sum^{n}_{i=1}\left(g(\alpha_{i})+\sum^{p}_{j=1}\left({\bm{\beta}}_{j}\frac{1}{n}\sum^{n}_{i=1}\Big(m(W_{i},(\partial g)\circ\phi_{j})-\alpha_{i}\widetilde{g}\left(X_{i},\phi_{j}(X_{i})\right)\Big)-|{\bm{\beta}}_{j}|\lambda\right)\right).$
From $\partial g\big(\alpha_{{\bm{\beta}}}(X_{i})\big)=\sum^{p}_{j=1}{\bm{\beta}}_{j}\widetilde{g}\left(X_{i},\phi_{j}(X_{i})\right)$, we have
$\max_{{\bm{\beta}}\in{\mathbb{R}}^{p}}\inf_{{\bm{\alpha}}\in{\mathcal{A}}^{n}}$ $\ \ \ \left\{\frac{1}{n}\sum^{n}_{i=1}\Big(g(\alpha_{i})-\alpha_{i}\partial g\big(\alpha_{{\bm{\beta}}}(X_{i})\big)+m\big(W_{i},(\partial g)\circ\alpha_{{\bm{\beta}}}\big)\Big)+\lambda\|{\bm{\beta}}\|_{1}\right\}.$
Consider the problem
$\inf_{{\bm{\alpha}}\in{\mathcal{A}}^{n}}\ \ \ \left\{\frac{1}{n}\sum^{n}_{i=1}\Big(g(\alpha_{i})-\alpha_{i}\partial g\big(\alpha_{{\bm{\beta}}}(X_{i})\big)+m\big(W_{i},(\partial g)\circ\alpha_{{\bm{\beta}}}\big)\Big)+\lambda\|{\bm{\beta}}\|_{1}\right\}.$
$\max_{{\bm{\beta}}\in{\mathbb{R}}^{p}}$ $\ \ \ \left\{\frac{1}{n}\sum^{n}_{i=1}\Big(g(\alpha_{{\bm{\beta}}}(X_{i}))-\alpha_{i}\partial g\big(\alpha_{{\bm{\beta}}}(X_{i})\big)+m\big(W_{i},(\partial g)\circ\alpha_{{\bm{\beta}}}\big)\Big)-\lambda\|{\bm{\beta}}\|_{1}\right\}.$
This is equivalent to
$\min_{{\bm{\beta}}\in{\mathbb{R}}^{p}}$ $\ \ \ \left\{\frac{1}{n}\sum^{n}_{i=1}\Big(-g(\alpha_{i})+\alpha_{i}\partial g\big(\alpha_{{\bm{\beta}}}(X_{i})\big)-m\big(W_{i},(\partial g)\circ\alpha_{{\bm{\beta}}}\big)\Big)+\lambda\|{\bm{\beta}}\|_{1}\right\}.$
Appendix C Preliminary for the Convergence Rate Analysis
@@ C.1 Rademacher complexity
C.2 Local Rademacher complexity bound
@@
Proposition C.1 (From Theorem 2.1 in Bartlett et al. (2005)).
@@ C.3 Bracketing entropy
C.4 Talagrand’s concentration inequality
@@
Proposition C.2 (Talagrand’s Lemma).
[…]
@@
Appendix D Basic inequalities
D.1 Strong convexity
Lemma D.1 ($L_{2}$ distance bound from Lemma 4 in Kato & Teshima (2021)).
If $\inf_{\alpha\in(-\infty),\infty}g^{\prime\prime}(\alpha)>0$, then there exists $\mu>0$ such that for all $\alpha\in{\mathcal{A}}$,
$\|\alpha-\alpha_{0}\|^{2}_{2}\leq\frac{2}{\mu}\Big(\text{BD}_{g}(\alpha)-\text{BD}_{g}(\alpha_{0})\Big)$
@@ D.1 Strong convexity
From the strong convexity and Lemma D.1, we have
@@
$\widehat{\alpha}\coloneqq\operatorname*{arg\,min}_{\alpha\in{\mathcal{A}}}\widehat{\text{BD}}_{g}(\alpha)+\lambda J(\alpha),$
@@ D.3 Risk bound
Then, from Proposition D.2, we have
@@
Proof of Proposition D.2 is trivial. We prove Proposition D.3 below.
@@
From the strong convexity and Lemma D.1, we have
@@
From Proposition D.2, we have
@@
Appendix E Proof of Theorem 6.1
@@
in Proposition D.3. We can bound this term by using the empirical-process arguments.
Note that since $\alpha_{0}\in{\mathcal{A}}$, it holds that $\alpha^{*}=\alpha_{0}$, which implies that
E.1 Preliminary
@@
Definition E.1 (Derived function class and bracketing entropy (from Definition 4 in Kato & Teshima (2021))).
Given a real-valued function class $\mathcal{F}$}, define $\ell\circ{\mathcal{F}}:=\{\ell\circ f\colon f\in{\mathcal{F}}\}$. By extension, we define $I:\ell\circ{\mathcal{A}}\to[1,\infty)$ by $I(\ell\circ h)=I(\alpha)$ and $\ell\circ{\mathcal{A}}_{M}:=\{\ell\circ\alpha\colon\alpha\in{\mathcal{A}}_{M}\}$. Note that, as a result, $\ell\circ{\mathcal{A}}_{M}$ coincides with $\{\ell\circ\alpha\in\ell\circ{\mathcal{A}}:I(\ell\circ h)\leq M\}$.
Proposition E.1.
@@
$H_{B}\big(\delta,\ell\circ{\mathcal{A}},\|\cdot\|_{L_{2}(P_{0})}\big)$ $\leq\frac{(s+1)(2v)^{\gamma}}{\gamma}\left(\frac{M}{\delta}\right)^{\gamma}.$
@@
$\sup_{\ell\circ\alpha\in\ell\circ{\mathcal{A}}_{M}}\|\ell\circ\alpha-\ell\circ\alpha^{*}\|_{L_{2}(P_{0})}$ $\leq c_{0}vM,$
$\sup_{\stackrel{{\scriptstyle\ell\circ\alpha\in\ell\circ{\mathcal{A}}_{M}}}{{\|\ell\circ\alpha-\ell\circ\alpha^{*}\|_{L_{2}(P_{0})}\leq\delta}}}\|\ell\circ\alpha-\ell\circ\alpha^{*}\|_{\infty}$ $\leq c_{0}vM,\quad\text{for all }\delta>0.$
Proposition E.2 (Lemma 5.13 in van de Geer (2000), Proposition 1 in Kanamori et al. (2012)).
@@
E.2 Upper bound using the empirical-process arguments
From Propositions E.1E.3, we obtain the following result.
Proposition E.4.
@@ E.2 Upper bound using the empirical-process arguments
$=O_{p}\left(\max\left\{\frac{\|\widehat{\alpha}-\alpha^{*}\|_{L^{2}(P_{0})}^{1-\gamma/2}\left(1+\left\|\widehat{\alpha}\right\|_{{\mathcal{A}}}\right)^{1+\gamma/2}}{\sqrt{n}},\frac{\left(1+\left\|\widehat{\alpha}\right\|_{{\mathcal{A}}}\right)^{2}}{n^{2/(2+\gamma)}}\right\}\right),$
@@
From Proposition D.3 and $\alpha_{0}\in\cal\alpha^{\text{RKHS}}$, we have
$\left\|\widehat{\alpha}(X)-\alpha_{0}(X)\right\|^{2}_{L_{2}(P_{0})}+\lambda\|\widehat{\alpha}\|^{2}_{{\mathcal{A}}}$
$\leq\left({\mathbb{E}}-\widehat{{\mathbb{E}}}\right)\left[L(\widehat{\alpha},D,X)-L(\alpha_{0},D,X)\right]+\lambda\|f_{0}\|^{2}_{{\mathcal{A}}}.$
From Proposition E.4, we have
$\left\|\widehat{\alpha}(X)-\alpha_{0}(X)\right\|^{2}_{L_{2}(P_{0})}+\lambda\|\widehat{f}\|^{2}_{{\mathcal{A}}}$
$=O_{p}\left(\max\left\{\frac{\|\widehat{\alpha}-\alpha_{0}\|_{L^{2}(P_{0})}^{1-\gamma/2}\left(1+\left\|\widehat{f}\right\|_{{\mathcal{A}}}\right)^{1+\gamma/2}}{\sqrt{n}},\frac{\left(1+\left\|\widehat{\alpha}\right\|_{{\mathcal{A}}}\right)^{2}}{n^{2/(2+\gamma)}}\right\}\right)+\lambda\|r_{0}\|^{2}_{{\mathcal{A}}}.$
@@
$\left\|\widehat{\alpha}(X)-\alpha_{0}(X)\right\|^{2}_{L_{2}(P_{0})}+\lambda\|\widehat{f}\|^{2}_{{\mathcal{A}}}=O_{p}(\lambda),$ (9)
$\left\|\widehat{\alpha}(X)-\alpha_{0}(X)\right\|^{2}_{L_{2}(P_{0})}+\lambda\|\widehat{f}\|^{2}_{{\mathcal{A}}}=O_{p}\left(\frac{\|\widehat{f}-f_{0}|_{L^{2}(P_{0})}^{1-\gamma/2}\left(1+\left\|\widehat{f}\right\|_{{\mathcal{A}}}\right)^{1+\gamma/2}}{\sqrt{n}}\right),$ (10)
$\left\|\widehat{\alpha}(X)-\alpha_{0}(X)\right\|^{2}_{L_{2}(P_{0})}+\lambda\|\widehat{f}\|^{2}_{{\mathcal{A}}}=O_{p}\left(\frac{\left(1+\left\|\widehat{f}\right\|_{{\mathcal{A}}}\right)^{2}}{n^{2/(2+\gamma)}}\right).$ (11)
@@
$\lambda\|\widehat{f}\|^{2}_{{\mathcal{A}}}=O_{p}(\lambda).$
Therefore, we have $\left\|\widehat{\alpha}(X)-\alpha_{0}(X)\right\|_{P_{0}}=O_{p}(\lambda^{1/2})$ and $\|\widehat{r}\|_{{\mathcal{A}}}=O_{p}(1)$.
@@
$\lambda\|\widehat{f}\|^{2}_{{\mathcal{A}}}=O_{p}\left(\frac{\|\widehat{f}-f_{0})\|_{L^{2}(P_{0})}^{1-\gamma/2}\left(1+\left\|\widehat{f}\right\|_{{\mathcal{F}}}\right)^{1+\gamma/2}}{\sqrt{n}}\right).$
@@
$\lambda\|\widehat{f}\|^{2}_{{\mathcal{A}}}$ $=O_{p}\left(\frac{\|\widehat{f}-f_{0})\|_{L^{2}(P_{0})}^{1-\gamma/2}\left(1+\left\|\widehat{f}\right\|_{{\mathcal{F}}}\right)^{1+\gamma/2}}{\sqrt{n}}\right)$
@@
$\|\widehat{f}\|_{{\mathcal{A}}}$ $=O_{p}\left(\frac{\left(1+\left\|\widehat{f}\right\|_{{\mathcal{F}}}\right)^{2}}{\lambda^{1/2}n^{2/(2+\gamma)}}\right)=o_{p}(1).$
@@
$\lambda\|\widehat{f}\|^{2}_{{\mathcal{A}}}=O_{p}\left(\frac{\left(1+\left\|\widehat{f}\right\|_{{\mathcal{F}}}\right)^{2}}{n^{2/(2+\gamma)}}\right).$
As well as the argument in (10), we have $\|\widehat{r}\|_{{\mathcal{A}}}=o_{p}(1)$. Therefore, we have
@@
Appendix F Proof of Theorem 6.2
@@
F.1 Proof of Lemma F.1
We prove Lemma F.1 by bounding (8) in Proposition D.3.
@@
Lemma F.2 (Corresponding to (26) in Zheng et al. (2022)).
Suppose that the conditions in Lemma F.1 hold. Then, for any $z>0$, with probability $1-\exp(-z)$ it holds that
@@
Lemma F.3 (Corresponding to (29) in Zheng et al. (2022)).
Suppose that the conditions in Lemma F.1 hold. If there exists $u_{0}>0$ such that
@@
Proposition F.4 (From (32) in Zheng et al. (2022)).
@@
Proposition F.5 (Corresponding to (36) in Step 3 of Zheng et al. (2022)).
Suppose that the conditions in Lemma F.1 hold. Then, there exists a universal constant $C>0$ such that
@@
Proposition F.6 (Upper bound of the Rademacher complexity).
Suppose that the conditions in Lemma F.1 hold. If $n\geq\text{Pdim}({\mathcal{F}}^{\text{FNN}})$, $u_{0}\geq 1/n$, and $n\geq(2eM)^{2}$, we have
@@
Then, we prove Lemma F.1 as follows:
Proof of Lemma F.1.
@@
then from (8) and Lemmas F.2 and F.3, for every $z>0$, there exists a constant $C>0$ independent $n$ such that
@@
If there exists $j\leq\ell$ such that $\widehat{f}\in\overline{{\mathcal{S}}}^{\text{FNN}}_{j}$, then from (F.1), for every $z>0$, with probability at least $1-8\exp(-z)$, there exists a constant $C>0$ independent of $n$ such that
@@
Here, to obtain (16), we used $\overline{u}\geq\max\left\{\sqrt{\log(n)/n},4\sqrt{3}M\sqrt{z/n},u^{\dagger}\right\}$, (F.1), (14), and (15).
From Proposition F.5, it holds that
@@
F.2 Proof of Lemma F.2
From Proposition C.1, we have
@@
This is a direct consequence of Proposition C.1. Note that $\alpha^{*}$ and $\alpha_{0}$ are fixed, and it is enough to apply the standard law of large numbers; that is, we do not have to consider the uniform law of large numbers. However, we can still apply Proposition C.1, which is a general than the standard law of large numbers, with ignoring the Rademacher complexity part.
@@
Then, from Proposition C.1, for every $z>0$, with probability at least $1-\exp(-z)$, it holds that